OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5). Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (anonymous):

For part 1, the first step would be to determine the slope of such a line. Simply evaluate \(\dfrac{\text{rise}}{\text{run}}\), or equivalently \(\dfrac{\Delta y}{\Delta x}\). Then apply point-slope form: The equation of a line that passes through \((x_0,y_0)\) with slope \(m\) is given by \(y-y_0=m(x-x_0)\). I'll let you figure out Part 2...give it a little thought.